The Harvard Family of Bibliography Styles

Peter Williams (peterw@archsci.arch.su.edu.au)

Thorsten Schnier (thorsten@archsci.arch.su.edu.au)

March 5, 1996

Contents

1 Introduction 1

2 Citations 1 2.1 Complete Citations . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2.2 Citation Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Partial Citations . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Exceptions for Individual Citations . . . . . . . . . . . . . . . . . 3

3 Styles 4 3.1 Bibliography Styles . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.2 Citation Styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.3 Parenthesis Style . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.4 Conjunction Style . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4 World Wide Web (WWW) References 5

5 Doing It By Hand 5

6 Acknowledgement 6

1 Introduction

This document describes the harvard family of bibliographic styles which are provided in addition to those described in Lamport (1986) and Patashnik (1988). This style is primarily intended for use with the BibTEX bibliographic database management system. However, provision is also made for hand coding of bibliographies.

2 Citations

2.1 Complete Citations

There are two primary forms of citation in the harvard style dependent upon whether the reference is used as a noun or parenthetically. Additionally, where there are more than two authors, all authors are listed in the first citation and in subsequent citations just the first author’s name followed by ‘et al.’ is used. The following example from Pitson (1978) illustrates these points.

The major improvement concerns the structure of the interview (Ulrich & Trumbo 1965, p. 112) …. Later reports (Carlson, Thayer, Mayfield & Peterson 1971) record greatly increased interviewer reliability for structured interviews. Wright (1969, p. 408) comments that ‘undoubtedly interviewer skill is directly related to the validity, quantity and quality of the interview output’, and this would suggest some sort of interviewer training is called for. Rowe (1960), for example, found that trained interviewers are better able to evaluate applicants with some measure of reliability. In addition Wexley, Sanders & Yukl (1973) showed that by extensive interviewer training all significant contrast effects could be eliminated. The results of the 1971 study (Carlson et al. 1971) are still relevant, but efforts to ….

To facilitate using a citation as a noun a new command \citeasnoun has been created which has the same syntax as the \cite command except that multiple citations are not permitted. The effect of this command is that

As \citeasnoun{btxdoc} and \citeasnoun[Annex~B]{latex} describe \ldots

produces

As Patashnik (1988) and Lamport (1986, Annex B) describe …

whereas

The \BibTeX\ \cite{btxdoc} and \LaTeX\ \cite[Annex~B]{latex} manuals \ldots

produces

The BibTEX (Patashnik 1988) and L^{A}TEX (Lamport 1986, Annex B)

manuals …

A second new command \possessivecite is provided for those instances where it is desired to use the citation as a possessive noun phrase. This is a variation on the \citeasnoun command and multiple citations are not permitted. As an example of its use

\possessivecite{latex} description of this feature is \ldots

produces

Lamport’s (1986) description of this feature is … A third new command \citeaffixed allows text to be affixed inside the beginning of the parenthesis of a parenthetical citation. This command is like the the \cite command except that it takes a second argument – the text to be affixed after the opening parenthesis. For example

\BibTeX\ manuals \citeaffixed{latex,btxdoc}{e.g.} describe \ldots

produces

BibTEX manuals (e.g. Lamport 1986, Patashnik 1988) describe …

2.2 Citation Modes

By default, where appropriate, citations are abbreviated automatically after the first reference when bibliographies are produced by BibTEX. Provision is also made for this feature to be accessed during manual coding. This feature may be overridden by using the \citationmode command which takes full, abbr or default as its single argument. The command \citationmode{full} makes the system use full citations, \citationmode{abbr} makes the system use abbreviated citations and \citationmode{default} causes the default behaviour of using full citations for the first instance and abbreviated citations thereafter.

Alternatively, the citation mode may be selected as an full, abbr or default option to the \usepackage command that invokes the harvard package. Use of the default option is redundant in that, if no citation mode option is used, that mode will be selected automatically.

2.3 Partial Citations

In addition to the primary forms of citation, the citation commands \citeyear and \citename are provided as building blocks for more complex citations that authors may (from time to time) require. \citeyear behaves like the \cite command except that only the year portion of the citation label is used. For example,

\citeyear{btxdoc,latex}

produces (1988, 1986). The parenthesis around the year list may be omitted by modifying the command name with a single asterisk (e.g. \citeyear*{btxdoc}). \citename behaves like the \citeasnoun command except that only the author name(s) portion of the citation label is used. For example,

\citename{btxdoc}

produces

Patashnik.

The use of these commands does not trigger the use of abbreviated citations for subsequent \citeasnoun and \cite references.

2.4 Exceptions for Individual Citations

Occasions arise where an author wishes to override the default behaviour for an individual citation (e.g. she may wish the citation to use the full list of authors where the default would use the abbreviated form). All commands that introduce authors’ names into a document (i.e. \cite, \citeasnoun, \citeaffixed, \possesivcite and \citename) may be modified with the addition of a single asterisk in order to force them to use the full list of authors names and by a double asterisk to force them to use the abbreviated form.

3 Styles

3.1 Bibliography Styles

There are six bibliography styles currently available within the harvard family, agsm (used in this document) which is based on Pitson (1978, pp. 95–98), dcu which is based upon the conventions in use in the Design Computing Unit, Department of Architectural and Design Science, University of Sydney, jmr for the Journal of Management Research, jphysicsB for the Journal of Physics B, kluwer which aspires to conform to the requirements of Kluwer Academic Publishers and nederlands which conforms to Dutch conventions. They are invoked by the \bibliographystyle as described in Lamport (1986, p. 74) and effect the layout of the entries in the bibliography.

Provided there is no name clash with other harvard options the bibliography style may be selected by passing it as an option to the \usepackage command that invokes the harvard package.

3.2 Citation Styles

There are two citation styles currently available within the harvard family, agsm (used in this document) and dcu which for the previous example would produce:

The BibTEX (Patashnik, 1988) and L^{A}TEX (Lamport, 1986, Annex B)

manuals …

and for multiple citations such as

  The original documentation \cite{btxdoc,latex} say \ldots

the agsm citation style produces

The original documentation (Patashnik 1988, Lamport 1986) say …

and the dcu citation style produces

The original documentation (Patashnik, 1988; Lamport, 1986) say …

The default citation style is agsm and both styles have no effect on the appearance of the \citeasnoun citation format.

These styles are invoked by the \citationstyle command, for example:

  \citationstyle{agsm}.

Because these styles affect the format of parenthetical citations, this command should appear before any \cite commands. Additionally the citation style may be selected by passing an option to the \usepackage command that invokes the harvard package. In order to avoid name clashes with the agsm and dcu bibliography styles the options agsmcite and dcucite are used with the \usepackage command in order to select agsm and dcu citation modes respectively.

3.3 Parenthesis Style

The type of parenthesis used in citations may be set using the \harvardparenthesis command which takes one argument. The argument to this command must be one of round, curly, angle, square or none. The default value is round. If it is a requirement that different parenthesis types are required for parenthetical cites that for the year portion for a \citeasnoun citation then the command \harvardyearparenthesis may be used to set the year parenthesis seperately. This command must be issued after any \harvardparenthesis command as that command sets both parenthetical and year parenthesis. If the bibliographic style chosen is agsm or dcu then the parenthesis style chosen using \harvardyearparenthesis is used with the year portion of the entries in the bibliographic listing. The options round, curly, angle, square and none may also be used with the the \usepackage command that invokes the harvard package.

Authors of style files for use with the harvard family who wish to make use of this feature should use the strings " \harvardleft " and " \harvardright " instead of the respective parenthesis characters where they wish them to be effected by the selection made with \harvardparenthesis.

3.4 Conjunction Style

In the previous examples for the agsm bibliographic style a “&” character is used to signify conjunction between a pair of names or between the last two names of a list of names. Similarly the word “and” is used for the dcu style. With these two styles this convention may be overwritten by using \renewcommand to redefine the command \harvardand. This should be done after the \citationstyle command (if used) as this command resets it to the default for the style selected.

4 World Wide Web (WWW) References

The agsm, dcu, jmr, jphysicsB and kluwer bibliographic styles support a new bibliographic entry field URL for specifying the URL of documents that are available via the World Wide Web. An example of this is the reference to Drakos’s (1994) documentation for his L^{A}TEX2html package in the file harvard.bib that is enclosed with the source for this document. When processed by L^{A}TEX2html documents using the harvard bibliographic package will have hypertext links created from the citation within the text to the reference list. If an entry in the reference list has an URL field then a hypertext link to the document will be created using the data in that field.

5 Doing It By Hand

Hand coding is accomplished much the same as described in Lamport (1986, p. 73) except that the new command \harvarditem is used in place of \bibitem. The syntax of this command is

\harvarditem [abbr-citation]{full-citation}{citation-year}{cite-key}

where

abbr-citation is the (optional) abbreviated citation (minus the year) to be used in the text subsequent to the first mention of a particular reference,

full-citation is the full citation (minus the year) to be used in the text on the first mention of a particular reference,

citation-year the year portion of the citation including any suffices required to disambiguate citations, and

cite-key is the key used in the \cite and \citeasnoun commands.

6 Acknowledgement

The motivation for this style came from Fay Sudweeks of the Design Computing Unit who also originated the formats for the dcu style and proofread their implementation.

The nederlands bibliographic style was implemented by Werenfried Spit (spit@vm.ci.uv.es).

The idea for \citeyear came from Renate Schmidt (Renate.Schmidt@mpisb.mpg.de).

The solution to the mysterious \enddocument problem came from Berwin A. Turlach (berwin@core.ucl.ac.be) as did the identification of a subtle problem with sorting entries in the reference list.

References

Drakos, N. (1994), The L^{A}TEX2html translator. Documentation accompanying the L^{A}TEX to html translator. *http://cbl.leeds.ac.uk/nikos/tex2html/doc/latex2html/latex2html.html

Lamport, L. (1986), L^{A}TEX: A Document Preparation System, Addison-Wesley.

Patashnik, O. (1988), BibTEXing. Documentation for general BibTEX users.

Pitson, J. (1978), Style Manual for authors editors and printers of Australian government publications, 3rd edn, Australian Government Publishing Service, Canberra.